OpenStudy (anonymous):

Medal reward Which of the following is the best estimate of 5 and three-eighths minus 1 and nine-tenths? 4 and one-half 2 and one-half 3 and one-half 4

OpenStudy (anonymous):

\[5\frac{3}{8} -1\frac{9}{10}=\frac{43}{8} -\frac{19}{10} =\frac{43\times10}{8\times10} -\frac{19\times8}{10\times8}\] \[=\frac{430}{80} -\frac{152}{80}=\frac{430-152}{80} =\frac{278}{80}=3.475\]

OpenStudy (anonymous):

Ans 3 and one-half

OpenStudy (anonymous):

\[3.475\approx 3.5 = 3 \frac{1}{2}\]